Jake Thomas began Thomas Refinishing Service on July 1, 2019. Selected accounts are shown below as of July 31, before any adjusting entries have been made.

Unadjusted Account Balances Debit Credit
Prepaid rent $5,700
Prepaid advertising 630
Supplies inventory 3,000
Performance obligation liability $600
Refinishing fees revenue 2,500


Using the following information, prepare the adjusting entries necessary on July 31 (b) in journal entry form.

1. On July 1, the firm paid one year’s advance rent of $5,700 in cash.
2. On July 1, $630 cash was paid to the local newspaper for an advertisement to run daily for the months of July, August, and September.
3. Supplies still available at July 31 total $1,100.
4. At July 31, refinishing services of $800 have been performed but not yet recorded or billed to customers. The firm uses the account Fees Receivable to reflect amounts due but not yet billed.
5. A customer paid $600 in advance for a refinishing project. At July 31, the project is one-half complete.

General Journal
Ref Description Debit Credit
1
To recordJuly rent expense.
2
To record July advertising expense.
3
To record supplies expense for July.
4
To record unbilled revenue earned in July.
5
To record advance fees earned in July.
Answer #1
Ref Description Debit Credit
1 Rent expense 475 =5700*1/12
    Prepaid rent 475
To record July rent expense.
2 Advertising expense 210 =630*1/3
      Prepaid advertising 210
To record July advertising expense.
3 Supplies expense 1900 =3000-1100
      Supplies inventory 1900
To record supplies expense for July.
4 Fees Receivable 800
       Refinishing fees revenue 800
To record unbilled revenue earned in July.
5 Performance obligation liability 300 =600*1/2
       Refinishing fees revenue 300
To record advance fees earned in July.